11. Determine the total resistance of the circuit illustrated to the right using the reciprocal formula introduced in class R1 2.2k03.3k0 R2 15 V 12. Construct the circuit illustrated to the right and, with the source 13. Use the total resistance to calculate the total circuit current. 14. Connect the source to the circuit and measure the total circuit current. 15. Determine the percent deviation of the resistance and current. 16. Record the measurements and calculations from steps 11 through 15 in Table 6 below. Table 6 Resistance Current Circuit Rr(MeasuredtI(calculated) IT(Measured) Variation Variation 17. Repeat steps 11 through 16 for the circuit illustrated to the R2 2.2k0 18. Repeat steps 11 through 16 for the circuit to the right using the equal value method to determine the total resistance. Assume nominal values to determine the calculated current15V and circuit resistance. 1k0 19. Determine the total power for each of the three circuits constructed above and record in table 3 (use measured values) Table 7 Pr Solve it and show the work for each circuits
Answer #1

11.

R R1 R2 2.2 3.3

\therefore R = 1.32 \Omega

13.

I=\frac{V}{R}=\frac{15}{1.32}=11.36 A

17.

\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}=\frac{1}{2.2}+\frac{1}{3.3}+\frac{1}{4.7}=0.97

\therefore R = 1.03 \Omega

I=\frac{V}{R}=\frac{15}{1.03}=14.56 A

18.

\frac{1}{R}=\frac{1}{R_1}+\frac{1}{R_2}+\frac{1}{R_3}=\frac{1}{1}+\frac{1}{1}+\frac{1}{1}=3

\therefore R = 0.33 \Omega

I=\frac{V}{R}=\frac{15}{0.33}=45.45 A